OpenStudy (anonymous):

Put the equation x^2+3x+7=9 in the form ax^2+bx+c=0

OpenStudy (harsimran_hs4):

you want on RHS so x^2+3x+7=9 x^2+3x+7-9 = 0 x^2+3x-2 = 0 a = 1, b = 3 c = -2

OpenStudy (anonymous):

Harsimran is right just subtract 9 on both sides